### v3

# autocorrelation function CFR channel frequency response

Abstract
This paper focuses on the parameterization of the multipath propagation model (MPM) for indoor broadband power line communications (PLC), which up to now has been established in an heuristic way. The MPM model was originally proposed for outdoor channels in the band up to 20 MHz, but its number of parameters becomes extremely large when used to model indoor channel frequency responses (CFR), which are much more frequency-selective than outdoor ones, and the band is extended to 80 MHz. This work proposes a fitting procedure that addresses this problem. It allows determining the model parameters that yield the best fit to each channel of a large database of single-input single-output (SISO) experimental measurements acquired in typical home premises of different European countries. Then, the properties of the MPM parameters are analyzed. The study unveils the relation between the model parameters and the main characteristics of the actual CFR like the frequency selectivity and the average attenuation. It also estimates the probability density function (PDF) of each parameter and proposes a fitting distribution for each of them. Moreover, the relationship among the main parameters of the model is also explored. Provided results can be helpful for the development of MPM-based models for indoor broadband PLC.
Index Terms:
Power line communications (PLC), channel model, multipath propagation model, channel frequency response, statistical characterization, probability density function.
I Introduction
The increasing demand of high-speed multimedia services led to reconsider the indoor power delivery network as a data transmission medium. The power line communication (PLC) technology is becoming an established solution since it exploits the existing wired infrastructure to convey high-speed data content, thus saving costs and deployment time. Furthermore, the latest standards and devices include advanced signal processing and enhanced transmission techniques, such as the exploitation of all the deployed network conductors and the bandwidth extension, that improve even further the achievable data rate, guaranteeing transmission rates in the order of gigabit per second at the physical layer [1, 2, 3]. This high performance level motivates a renewed interest in developing novel and updated PLC solutions, especially for home networking and smart grid applications [4, 5, 6, 7, 8].
In order to support the growth of PLC technology, it is fundamental to provide effective channel models, which are able to numerically emulate the properties of actual PLC networks avoiding on field tests. Two basic approaches can be tackled in defining a model, i.e., bottom-up and top-down, and both of them can follow a deterministic or a statistical strategy [9, 10, 11, 12]. A deterministic strategy typically considers a specific network configuration, while a statistical strategy attempts to introduce some variability by means of statistical distributions. Concerning the possible approaches, the bottom-up method has a strong physical connection and describes the phenomena involved in the signal propagation by exploiting the transmission line theory. Typically, it provides a faithful channel representation, although it turns out to be a computationally onerous procedure. Some examples concerning the single-input-single-output (SISO) channel modeling can be found in [13, 14], which propose a deterministic approach, and in [15, 16, 17, 18, 11], which follow a statistical approach. Recently, the models discussed in [19] and [20] extends the SISO bottom-up strategy to the multiple-input multiple-output (MIMO) case.
Conversely, the top-down approach is an empirical method that fits an analytic function to a measurements database. This strategy facilitates the statistical extension by adopting a certain distribution for the parameters that describe the theoretical function. One of the first and more common SISO models is the multipath propagation model (MPM) proposed in [9] and was later statistically extended in [21]. Extensions of the SISO formulation to the MIMO are presented in [22] and [23].
The PLC channel response also exhibits a short-term variation synchronous with the mains voltage, which allows modelling it as a linear periodically time-varying (LPTV) system [24]. However, since the time variation is quite small compated to the duration of the impulse response, i.e., the channel response is underspread, the slow variation approximation can be made and the channel response can be modeled as a series of time-invariant responses that repeat with the period of the mains signal. Hence, the MPM model can be applied to each of these time-invariant responses.
I-A Contribution
The top-down MPM model formulated in [9] has been widely used in the PLC scenario [25, 10, 26, 27, 28]. However, it was proposed for outdoor PLC channels in the band up to $20$ MHz. Its statistical extension to indoor channels made in [29] and [21] was based on heuristic approaches and no rigorous measurement-based analysis of this model for indoor scenarios in the band up to 80 MHz has been found in the literature.
In this context, we make three main contributions:
  • We propose a method to determine the parameters of the MPM model for indoor broadband PLC channels in the band up to $80$ MHz. It addresses the problem of the extremely large number of parameters that results when the MPM model in [9] is extended to indoor scenarios and that are a consequence of the much more frequency selective behavior of indoor channel frequency responses and of the bandwidth extension to $80$ MHz. Since interference with commercial frequency modulation (FM) broadcasting is important above $87$ MHz in many regions.
  • The proposed method has been used to perform a channel-by-channel fitting, determining the set of MPM parameters that yields the best approximation to each of the channels of a large database consisting of 426 CFRs measured in indoor scenarios of different European countries.
  • We give an in-depth analysis of the resulting MPM parameters, showing their relation to the main channel characteristics like the frequency selectivity and the average attenuation, assessing their statistical behavior and the relation between them. Unlike the a posteriori parameters inference discussed in previous works, this study unveils the intrinsic properties of the MPM parameter and can be helpful for the development of novel MPM models or the refinement of existent ones.
I-B Organization
The paper is organized as follows. First, the MPM is briefly recalled in Section II. Second, the assumptions and the method proposed to determine the MPM parameters that give the best fit to each measured channel is detailed and justified in Section III. The properties of the computed parameters, such as their relation to the channel characteristics, statistical behavior and relation between them are discussed in Section IV. Finally, conclusions follow in Section V.
II Background: The Multipath Propagation Model
The MPM follows a top-down approach and describes the multipath propagation of the power signal into the PLC medium that is due to the line discontinuities, such as branches, or unmatched loads. The analytic formulation of the CFR can be expressed as [9]
$H\left(f\right)=A\sum_{i=0}^{N-1}{g_{i}}e^{-(a_{0}+a_{1}f^{K})d_{i}}e^{-j2\pi fd_{i}/v},$ (1)
where $a_{0}$ and $a_{1}$ are the attenuation coefficients, $g_{i}$ and $d_{i}$ are the path gain and the path length of the $i$-th path [^1] For the sake of notation simplicity in the subsequent fitting process, paths have been indexed as $10\leq i\leq N-$ 1, instead of $1\leq i\leq N$ as done in [9]., respectively. The coefficient $g_{i}$ can be interpreted as the product of transmission and reflection coefficients by restricting $|g_{i}|\leq 1$ and by introducing the normalization coefficient $A>0$, which allows adjusting the attenuation of the resulting CFR [9]. The term $N$ denotes the number of paths and $K$ is the exponent of the attenuation factor related to the cable features. This work uses $N=2554$ (its rationale will be justified in Section III.III-B), which is much larger than the values used in outdoor channels. Typical values of $K$ range between 0.5 and 1 [9], hence $K=1$ is assumed for the rest of this paper. The parameter $\nu=c\epsilon_{r}$ is the propagation speed of light in the cables, where $c$ is the speed of light in the vacuum and $\epsilon_{r}$ is the relative dielectric constant of the insulator which envelops the conductors, which has been set to $\epsilon_{r}=1.5$ in order to take into account the non-uniformity of the dielectric [9].
III Determination of the MPM Parameters
In order to provide a statistical characterization of the parameters that describe the MPM formulation in (1), the first step is to compute the MPM parameters that give the best fitting of this analytic expression to each of the channels of the measurement database. The latter consists of $426$ indoor CFRs measured in different European countries using a vector network analyzer. CFRs have $M=1262$ frequency samples given by $f_{m}=f_{0}+m\Delta f$, with $f_{0}=1.0$ MHz, $10\leq m\leq M-$ 1 and $\Delta f=62.5978$ kHz. Part of this collection was previously used in [23] and [12], and the results are consistent with the ones obtained from other extensive measurements in Italy that were characterized in [30].
Since the model is described by many parameters, the following assumptions must be adopted in order to simplify the fitting strategy. First, the attenuation coefficients $a_{0}$ and $a_{1}$ are estimated through the robust regression fit of each channel measurement. This linear attenuation profile can assumed to be essentially determined by the longest path of the MPM. Hence, we consider the model in (1) with a single path of the largest length, $L$ (its value will be given in Section III.III-B). Since $A$ is a normalization factor used to force $|g_{i}|\leq 1$, when a single path is considered it happens that $|A\cdot g_{i}|=1$. Accordingly, the robust regression fit of the longest path in (1) can be expressed as
$10\log_{10}|H(f)|^{2}=-2(a_{0}+a_{1}f)L\cdot 10\log_{10}e=\alpha_{0}+\alpha_{1}f,$ (2)
where $\alpha_{0}$ and $\alpha_{1}$ are the coefficients of the robust regression fit. Thus, the attenuation coefficients are computed as
$a_{0}=-\frac{\alpha_{0}}{20L\log_{10}e},\quad a_{1}=-\frac{\alpha_{1}}{20L\log_{10}e}.$ (3)
Second, the path lengths are discretized yielding uniformly spaced distances,
$d_{i}=i\frac{L}{N}\quad\text{with }i=0,...,N-1.$ (4)
This is a viable assumption since indoor PLC networks typically exhibit a complex topology with a great number of branches. This results into approximately uniformly distributed distances (spaced $1.25$ m for the values of $L$ and $N$ considered in this work) traveled by the signal and its reflections from the transmitter side towards the receiver outlet.
Afterwards, the path gains, $g_{i}$ that give the best fitting to each each measured channel have to be determined. This could be done by means of a LS estimation that minimizes the root mean square error (RMSE) w.r.t. the measured channel in the set of $M$ measured frequencies $f_{m}$,
$\text{RMSE}=\sqrt{\frac{1}{M}\sum_{m=0}^{M-1}{|H(f_{m})-\hat{H}(f_{m})|^{2}}},$ (5)
where $H(f_{m})$ and $\widehat{H}(f_{m})$ represent the measured and fitted CFR at frequency $f_{m}$, respectively.
However, the RMSE is dominated by the error in the largest values of the CFR. Since the amplitude response of PLC is strongly frequency-selective, this causes the fitting to be very accurate at lowly attenuated values to the detriment of the fitting of the highly attenuated values. To avoid this end and achieve a balanced error distribution along the entire amplitude range, we determine the path gains that minimize the normalized root mean square error (NRMSE) defined as
$\text{NRMSE}=\sqrt{\frac{1}{M}\sum_{m=0}^{M-1}{\frac{|H(f_{m})-\hat{H}(f_{m})|^{2}}{|H(f_{m})|^{2}}}}.$ (6)
The $g_{i}$ can be obtained by solving a weighted (WLS) problem formulated by expressing (1) in matrix form as follows,
$\underbrace{\begin{bmatrix}H(f_{0})\\ \vdots\\ H(f_{M-1})\\ H^{*}(f_{M-1})\\ \vdots\\ H^{*}(f_{0})\\ \end{bmatrix}}_{\mathbf{h}}=\underbrace{\begin{bmatrix}P_{0,0}&\dots&P_{0,N-1}\\ \vdots&&\vdots\\ P_{M-1,0}&\dots&P_{M-1,N-1}\\ P^{*}_{M-1,0}&\dots&P^{*}_{M-1,N-1}\\ \vdots&&\vdots\\ P^{*}_{0,0}&\dots&P^{*}_{0,N-1}\\ \end{bmatrix}}_{\mathbf{P}}\underbrace{\begin{bmatrix}g_{0}\\ \vdots\\ g_{N-1}\end{bmatrix}}_{\mathbf{g}}$ (7)
where $P_{m,i}=\exp(-(a_{0}+a_{1}f_{m})d_{i})\cdot\exp(-j2\pi f_{m}d_{i}/v)$ and the scaling factor $A$ has been set to 1 by now.
By defining the weighting matrix
$\begin{split}\mathbf{W}=\text{diag}\bigg{\{}&\frac{1}{|H(f_{0})|^{2}},\frac{1}{|H(f_{1})|^{2}},\ldots,\frac{1}{|H(f_{M-1})|^{2}},\\ &\frac{1}{|H(f_{M-1}|^{2}},\ldots,\frac{1}{|H(f_{1})|^{2}},\frac{1}{|H(f_{0})|^{2}}\bigg{\}},\end{split}$ (8)
the values of $g_{i}$ can be obtained as
$\mathbf{g}=\left(\mathbf{P}^{\textrm{H}}\mathbf{W}\mathbf{P}\right)^{-1}\mathbf{P}^{\textrm{H}}\mathbf{W}\mathbf{h}.$ (9)
III-A Dominant Paths Selection Procedure
In order to fit each measured channel with a reduced set of parameters, paths with negligible contribution can be deleted, while keeping the most significant ones. To this end, a decimation procedure that iteratively discards the path with the smallest contribution to approximate the measured CFR while keeping the NRMSE below a certain threshold is employed. The index of the path to be discarded at each iteration, designated as $j$, is determined as
$j=\operatorname*{arg\,min}_{i}{\left\{\sum_{m}|g_{i}|e^{-(a_{0}+a_{1}f_{m})d_{i}}\right\}}.$ (10)
The rationale for this criterion is that the actual contribution of a path to the overall CFR is essentially determined by the factor $|g_{i}|\cdot\exp(-(a_{0}+a_{1}f)d_{i}$ in the MPM in (1).
The decimation procedure continues while the NRMSE (dB) remains lower than a given threshold, which has been set to $-20$ dB. This ensures that the channel computed with a reduced set of paths still represents a very good approximation to the corresponding measured channel. Since the columns of the matrix $\mathbf{P}$ are not orthogonal, the path gains $g_{i}$ have to be recomputed after a path is discarded. To this end, the problem in (7) is reformulated by removing the $j$-th column of the matrix $\mathbf{P}$ and the new path gains are obtained. Table I summarizes the presented fitting and decimation procedure applied to each measured channel, where the expressions to obtain the values of $L$ and $N$ will be given in Section III.III-B.
When the decimation procedure ends, the path gains obtained for the considered measured channel are normalized by their maximum absolute value. This normalization factor provides the value for the parameter $A$. This way, the $g_{i}$ values, which represent the product of transmission and reflection coefficients, are in the range $g_{i}\in[-1,1]$, and the normalization coefficient $A\in(0,\infty)$.
Table I: Algorithm of the fitting and path selection procedure applied to each measured channel
Algorithm: MPM fitting procedure
Input: $\mathbf{h}$ (measured CFR) and $\text{NRMSE}_{\text{threshold}}=-20$ dB
Output: $N$ (number of dominant paths), $\mathbf{d}$ (path lengths), $\mathbf{g}$ (path gains)
and $\mathbf{\widehat{h}}$ (fitted version of $\mathbf{h}$)
1: Initialize $L$ and $N$ using (12) and (15), respectively
2: Initialize $\mathbf{d}$ using (4)
3: Compute $a_{0}$ and $a_{1}$ using (3)
4: Compute $\mathbf{g}$ using (9)
5: Compute $\mathbf{\widehat{h}}=\mathbf{P}\mathbf{g}$
6: Compute the NRMSE using (6)
7: while NRMSE (dB)$<\text{NRMSE}_{\text{threshold}}$ do
     8: Obtain $j$ using (10)
     9: Calculate $\mathbf{P}=[\mathbf{P}(:,0:j-1),\mathbf{P}(:,j+1:N-1)]$
     10: Update $\mathbf{g}$ using (9)
     11: Update $\mathbf{\widehat{h}}=\mathbf{P}\mathbf{g}$
     12: Compute the NRMSE using (6)
     13: Update $N=N-1$
14: end while
15: Compute $\mathbf{P}=[\mathbf{P}(:,0:j-1),\mathbf{P}(:,j),\mathbf{P}(:,j+1:N-1)]$
16: Update $\mathbf{g}$ using (9)
17: Update $\mathbf{\widehat{h}}=\mathbf{P}\mathbf{g}$
The number of initial paths considered for the model at the pre-decimation point is $N=2554$ (as justified in the next subsection). This gives average and standard deviation values of the NRMSE (computed over the set of measured channels) equal to $\overline{\textrm{NRMSE}}=-36.79$ dB and $\sigma_{\textrm{NRMSE}}=7.99$ dB, respectively. After the application of the proposed decimation procedure, the number of paths is reduced to an average value (over the whole set of fitted channels) of 217.84, i.e., less than 10% of the initial value, and the corresponding average and standard deviation values of the NRMSE (dB) are $\overline{\textrm{NRMSE}}=-20.05$ dB and $\sigma_{\textrm{NRMSE}}=0.0678$ dB, respectively. Hence, the proposed decimation process achieves a good trade-off between modeling complexity and accuracy.
For illustrative purposes, Fig.1 shows the amplitude and unwrapped phase of a measured CFR, $H(f)$, and its corresponding fitted response, $\hat{H}(f)$, obtained with $N=183$ dominant paths determined using the proposed decimation process. As can be observed, the measured trace is almost veiled by the estimated one, which highlights the notable accuracy of the fitting process.
The procedure described up to now assumed an initial number of paths, $N$, and a maximum path length, $L$. These parameters cannot be chosen arbitrarily and their selection criteria are described in the following section.
Figure 1: Amplitude (a) and unwrapped phase (b) response of a representative measured channel and its corresponding fitting determined with $N=183$ dominant paths obtained with the proposed decimation procedure.
III-B Relation Among L, M, and N
The maximum length $L$, the maximum number of paths $N$, and the number of frequency samples $M$ depend among each other and must satisfy certain constraints. Since only a discrete and finite number of frequencies is available, the expression in (1) given for the continuous frequency domain can be rewritten for the discrete domain as
$H(f_{m})=A\sum_{i=0}^{N-1}r_{i}e^{-j2\pi m\Delta fi\frac{L}{N\nu}},$ (11)
where $\Delta f=(f_{M-1}-f_{0})/(M-1)$ is the frequency resolution, while $\tau=L/(N\nu)$ is the time resolution. Note that the real part of the exponential term has been embedded within the corresponding path gain in a unique variable named $r_{i}$. The adopted notation does not affect the analysis since the real exponential part acts only as an attenuation term that increases with frequency. This way, equation (11) can be seen as a discrete time Fourier transform (DFT) with frequency index $m$ and time index $i$. As known, the DFT assumes the time signal belonging to the domain of integer multiples of $\tau$ and periodic of $N\tau$. Such a domain is denoted with $\mathbb{Z}(\tau)/\mathbb{Z}(N\tau)$. Hence, in the frequency domain the signal is defined in $\mathbb{Z}(\Delta f)/\mathbb{Z}(N\Delta f)$, with $\Delta f=1/(N\tau)$. Therefore, the relation between $L$ an the frequency samples $M$ is given by
$\Delta f=\frac{f_{M-1}-f_{0}}{M-1}=\frac{1}{N\frac{L}{N\nu}}=\frac{\nu}{L}\Rightarrow L=\frac{(M-1)\nu}{f_{M-1}-f_{0}}$ (12)
Since $f_{M-1}=79.9358$ MHz and $f_{0}=1$ MHz, $M=1262$ and the propagation speed is $\nu=2\cdot 10^{8}$m/s, the maximum possible length that can be assumed is $L=3195$ m. The limit of more than 3 kilometers turns out to be a maximum length that is long enough to describe any generic in-home network topology. Moreover, always referring to the DFT representation of the CFR reported in (11), the DFT frequency period, i.e., $N\Delta f$, must be greater or equal than two times the maximum channel frequency, i.e., $f_{M-1}$, in order to avoid aliasing problems according to the sampling theorem. Thus, the relationship among $N$ and $L$ turns out to be
$N\Delta f=\frac{1}{\tau}\geq 2f_{M-1}\Rightarrow N\geq\frac{2f_{M-1}L}{\nu}.$ (13)
If the aforesaid values for $f_{M-1}$ and $\nu$ are considered, the relation in (13) becomes $N\geq 0.8L$. A final relation among $N$ and $L$ can be noticed by rewriting the imaginary part of the exponential term in (1) as
$e^{-j2\pi fd_{i}/\nu}=e^{-j2\pi kd_{i}/\lambda}$ (14)
where $\lambda=\nu/\Delta f$ is the maximum possible wavelength, which is related to the frequency resolution $\Delta f$. The minimum observable wavelength is $\lambda_{min}=\nu/f_{M-1}$. Hence, to avoid numerical errors, the shortest (non-zero) path length $d_{1}=\frac{L}{N}$ must be greater than $\lambda_{min}/2$, according to the sampling theorem. This provides the last relationship, expressed by
$d_{1}=\frac{L}{N}\geq\frac{\lambda_{min}}{2}\Rightarrow N\leq\frac{2f_{M-1}L}{\nu}$ (15)
It can be noted that, given the above mentioned specifications, as well as the relations (13) and (15), the relationship among $L$ and $N$ must hold with equality, i.e., $N=0.8L=2554$ for the considered database.
IV Analysis of the MPM Parameters
This section analyzes the statistical properties of the MPM parameters obtained according to the fitting procedure detailed in Section III for the overall set of measured channels. First, the relation of some model parameters to the main magnitudes used to characterize the channel response, such as the delay spread and the average channel gain. Then, the PDF of the MPM parameters is estimated and, finally, the statistical relation between the model parameters is studied.
IV-A Relation to the Channel Characteristics
The parameters $a_{0}$ and $a_{1}$ where obtained from the robust regression fit of the amplitude response of the measured channels, with $a_{1}$ being related to the slope of the attenuation profile and $a_{0}$ with the y-intercept. Hence, the latter parameter has to be somehow related to the average channel gain, defined as $G\textrm{ (dB)}=\frac{1}{M}\sum_{m=0}^{M-1}20\log_{10}{|H(f_{m})|}$. This relation can be clearly observed in Fig. 2 (a), where the scatter plot of $a_{0}$ vs $G$ (dB) is depicted, along with its robust fit regression line, whose expression is given in Table II. As expected, channels with larger average gains need lower values of $a_{0}$ to be synthesized.
Figure 2: Scatter plot of (a) the value of $\bm{a_{0}}$ vs the average channel gain of each measured channel, $\bm{G}$ (dB); (b) the number of dominant paths vs the average gain and the delay spread of the channels; (c) value of $\bm{A}$ vs the average channel gain of each measured channel.
The number of dominant paths needed to fit a given CFR is related to the frequency selectivity of the channel, e.g., a single path suffices to synthesize a flat channel. The frequency selectivity of a channel is usually quantified by means of the coherence bandwidth or its inversely related magnitude, the delay spread (also referred to as multipath spread) [31]. In PLC these magnitudes are known to be related to the average channel gain [8]: channels with lower average channel gains tend to have larger delay spreads. Hence, these magnitudes are expected to be related to the number of dominant paths, $N$. Fig. 2 (b) illustrates this relation by showing the scatter plot of the number of dominant paths vs the average channel gain and the delay spread of the corresponding measured channel. Its robust regression plane, whose expression is given in Table II, is also depicted for reference. The goodness of this fit is confirmed by the NRMSE between the actual and predicted values of $N$, which is $-22$ dB.
The value of the normalization parameter $A$ is also related to the average channel gain, as shown in Fig. 2 (c), which depicts the scatter plot of $A$ vs $G$ (dB) and its corresponding exponential fitting function. The parameters of the latter, which have been obtained by means of a LS fitting, are given in Table II. As observed, the lower the average channel gain, the lower the value of $A$ needed to synthesize the channel.
Table II: Parameters of the fitting expressions displayed in Fig. 2.
Fitting expression $\bm{\alpha}$ $\bm{\beta}$ $\bm{\gamma}$
[…]
$\widehat{A}=\alpha\cdot e^{\beta\cdot G\textrm{ (dB)}}$ $8.2517\cdot 10^{-1}$ $1.0636\cdot 10^{-1}$
IV-B Statistical Characterization
This subsection estimates the chief moments and the PDF of the MPM parameters, computed for the overall set of measured channels, that have been obtained using the fitting procedure presented in Section IV. For each empirical (estimated) PDF, the probability distribution that gives the best fit is given. To this end, the most common distribution functions are considered: beta, Birnbaum-Saunders, exponential, gamma, generalized extreme value (GEV), Gumbel, inverse Gaussian, logistic, log-logistic, log-normal, Nakagami, normal, Poisson, Rayleigh, Rician, $t$ location-scale and Weibull. The parameters of each distribution that yield the best fit to the set of MPM parameters are obtained by means of the maximum likelihood (ML) estimation assuming independent samples [32]. Afterwards, the Anderson-Darling goodness-of-fit test is applied and the distribution yielding the highest $p$-value is selected. When the test gives the same $p$-value for all distributions or the null hypothesis is rejected for all of them, the one with the highest value of the likelihood function is selected. For some parameters, more elaborate models consisting of mixture distributions have been considered.
The mean and standard deviation of the MPM parameters are given in Table III. The fitting statistical distributions with their corresponding parameter values are summarized in Table IV. These reported results allow future replicability of this study and simplify the development of improved statistical models.
Table III: Mean and standard deviation of the MPM parameters
MPM parameter mean standard deviation
Atten. coeff. ($a_{0}$) 0.0011 5.0089e-4
Atten. coeff. ($a_{1}$) 6.0953e-12 4.1137e-12
Norm. coeff. ($A$) 0.0422 0.0649
Number of dominant paths ($N$) 217.8427 1.3457
Path gains ($|g_{i}|$) 0.1384 0.1859
Path lengths ($d_{i}$) 343.9491 633.4392
Figure 3: Empirical and fitted PDFs of the attenuation coefficients $\bm{a_{0}}$ and $\bm{a_{1}}$ and the normalization coefficient $\bm{A}$.
Figure 4: Empirical and fitted PDFs of the number of dominant paths, $\bm{N}$, the modulus of the path gains $\bm{|g_{i}|}$ and the path lengths $\bm{d_{i}}$.
The empirical s of the MPM parameters are reported in Fig. 3 and Fig. 4. Regarding $a_{0}$ and $a_{1}$, whose EPDFs are show in Fig. 3 (a) and (b), respectively, the distribution that gives the best fit is the GEV ($p$-value$=0.286$ for $a_{0}$ and 0.8941 for $a_{1}$), although the normal distribution also gives a high $p$-value (0.1967 for $a_{0}$ and 0.5664 for $a_{1}$). In practice, the GEV is a family of three types of extreme value distributions, namely type I, or Gumbel (or log-Weibull) [33], type II, or Fréchet, and type III, or Weibull distribution. This result makes sense since the GEV is the limit distribution of properly normalized maxima of a sequence of independent and identically distributed random variables. The GEV distribution for $a_{0}$ is justified since it represents the maximum (or the minimum in case of negative values of $a_{1}$) of the channel attenuation profile for the identified trend of each measured CFR.
Regarding the normalization coefficient, $A$, whose EPDF is displayed in Fig. 3 (c), the Log-logistic gives the largest $p$-value (0.909) although, although the GEV and the log-normal also give high values ($p$-values $0.788$ and $0.4536$, respectively). The latter is not surprising, as $A$ is determined as the maximum value of the path gains, $|g_{i}|$, of each fitted channel.
Regarding the probability distribution of the number of dominant paths, $N$, while this magnitude can only take non-negative integer values, discrete distributions like the binomial and Poisson ones give very poor fitting to the actual set of values. Hence, $N$ has been considered as a discretized version of a continuous random variable. This same approach has been also applied to fit the path length, $d_{i}$. By doing so, the GEV gives the best fit ($p$-value$=0.86752$) to the number of dominant paths, $N$, whose EPDF is displayed in Fig. 4 (a). As a matter of fact, $N$ represents the maximum number of dominant paths that actively contribute to reconstruct each measured CFR.
Concerning the characterization of the path gain, its absolute value, $|g_{i}|$, is considered for the sake of simplicity, as the EPDF of $g_{i}$ shows a symmetric behavior around the $y$-axis. The statistical characterization of $|g_{i}|$ must take into account that the normalization process used to force $g_{i}\in[-1,1]$ in each fitted channel causes the probability of $|g_{i}|=1$ to be non-zero. Hence, $|g_{i}|$ has to be modeled by a mixed random variable (RV) consisting of a continuous component and a discrete term at $|g_{i}|=1$. Fig. 4 (b) shows that the continuous term of the EPDF can be fitted by a log-normal RV. Accordingly, the overall PDF can be expressed as $f(|g_{i}|)=\pi_{0}q(|g_{i}|\big{|}\mu,\sigma)+\pi_{1}\delta(|g_{i}|-1)$, where $q(\cdot\big{|}\mu,\sigma)$ denotes the PDF of a log-normal RV with parameters $\mu$ and $\sigma$ and $\delta(\cdot)$ is the Dirac delta function. The value of these parameters are given in Table IV.
The EPDF of $d_{i}$, displayed in Fig. 4 (c), exhibits a bimodal behavior, with maxima around $d_{i}=50$ m and $d_{i}=3195$ m. The latter is an effect of the use of the discrete frequency domain version of (1), which obliges to set a limit for the largest path length. Since the observed bimodal behavior cannot be obtained with any of the distributions indicated above, we have resorted to mixture models. It has been found that the Weibull distribution gives a good fitting to the values of $d_{i}\leq 1500$ while the GEV, shifted and reflected about $d_{N-1}=\frac{N-1}{N}L$, adequately fits the values of $d_{i}>1500$. Hence, the PDF of $d_{i}$ can be modeled as $f(d_{i})=\pi_{0}g_{0}(d_{i}\big{|}\lambda_{0},k_{0})+\pi_{1}g_{1}(d_{N-1}-d_{i}\big{|}k_{1},\sigma_{1},\mu_{1})$, where $g_{0}(\cdot\big{|}\lambda_{0},k_{0})$ denotes the PDF of a Weibull distribution with scale and shape parameters $\lambda_{0}$ and $k_{0}$, respectively, and $g_{1}(\cdot\big{|}k_{1},\sigma_{1},\mu_{1})$ is the PDF of a GEV distribution with location, scale and shape parameters $k_{1}$, $\sigma_{1}$ and $\mu_{1}$, respectively, whose values are given in Table IV. As shown in Fig. 4 (c), the PDF of the proposed mixture gives a good fitting to the empirical one.
Table IV: Fitting distributions of the MPM parameters
MPM
parameter
Distribution
Distribution
parameters
Atten. coeff.
$a_{0}$
GEV
$k=-0.2744$
$\sigma=4.9492\cdot 10^{-4}$
$\mu=9.3370\cdot 10^{-4}$
Atten. coeff.
$a_{1}$
GEV
$k=-0.1781$
$\sigma=3.8980\cdot 10^{-12}$
$\mu=4.4536\cdot 10^{-12}$
Norm. coeff.
$A$
Log-logistic
$\mu=-3.8584$
$\sigma=0.6710$
Number of
dominant paths
$N$
GEV
$k=0.1641$
$\sigma=84.3237$
$\mu=153.5548$
Path gains
$|g_{i}|$
Mixture
(lognormal, $|g_{i}|=1$)
$\pi_{0}=1-\pi_{1}$
$\pi_{1}=4.6017\cdot 10^{-3}$
$\mu=-2.9139$
$\sigma=1.5445$
Path lengths
$d_{i}$
Mixture
(Weibull, GEV)
$\pi_{0}=9.5074\cdot 10^{-1}$
$\pi_{1}=1-\pi_{0}$
$\lambda_{0}=2.1894\cdot 10^{2}$
$k_{0}=1.2314$
$k_{1}=1.3432$
$\sigma_{1}=42.5933$
$\mu_{1}=27.4299$
In order to get an in-depth analysis of the behavior of $d_{i}$, Fig. 5 (a) displays the path length, $d_{i}$, vs the dominant path index, $i$, for some illustrative measured channels. For the sake of clarity, the markers corresponding the same channel have been linked by a dashed line of the the same color. As seen, all channels have a set of paths whose length is uniformly distributed from zero (or almost zero) and a given value that depends on the channel. However, some channels, like Ch. 3-7, also have a set of paths with very large length. In many cases, like Ch. 3 and Ch. 4, there are no paths with intermediate distances. This behavior is related to the frequency selectivity of the channel: the larger the frequency selectivity, the large the distances of the paths required to synthesize it. This can be observed in Fig. 5 (b), where the scatter plot of the median value of the dominant path lengths vs the delay spread of each measured channel is given. As observed, channels with higher delay spread values (larger frequency selectivity) need larger paths to be synthesized. Futhermore, it is also expected that the higher the number of dominant paths, the larger the median length of the paths. Fig. 5 (c) confirms this end by showing a strong linear relationship between them.
Figure 5: (a) Dominant path length vs dominant path index for some illustrative measured channels; (b) scatter plot of the median value of the dominant path length vs the delay spread for the set of measured channels; (c) scatter plot of the median value of the dominant path length vs the number of dominant paths for the set of measured channels.
IV-C Comparison with Other Works
The MPM parameter distributions that have been just presented are herein compared with those heuristically or empirically obtained in previous scientific works. For example the presented statistical analysis confirms the assumptions made in [21], which statistically extended the MPM in (1) relying on the initial idea presented in [10]. Hence, it was inferred in [21] that the path gains $g_{i}$ turn out to be log-normally distributed variables multiplied by a random flip sign. The main difference is found for the statistics of $N$, supposed to have a Poisson distribution in [21], but that results into a GEV one. This difference was already hypothesized in [21] and is also justified by the different considered databases. The path lengths, $d_{i}$, were assumed to be log-normally distributed in [21], while the mixture of a Weibull and GEV has been obtained in this work. Additionally, the characterization method proposed in this paper provides a statistics for the other parameters, i.e., $a_{0}$, $a_{1}$, and $A$.
Other results concerning the MPM parameters distribution are listed in [34](Tab. 1), that extends the model in (1) to the MIMO transmission. In this case, the path gains $g_{i}$ are assumed uniformly distributed in $[-1,1]$ and the attenuation coefficient $a_{1}$ turns out to be constant. While, $K$ is normally distributed and $a_{0}$ exhibits a shifted exponential distribution. Instead, the channel median $A$ is uniformly or exponentially distributed depending on the considered channels. The same results are also summarized in [35, Tab. 2.10].
IV-D Relationship Between Parameters
The development of MPM-based models requires a complete characterization of the involved parameters. This includes both their marginal PDFs, which have given in the previous subsection, and their joint distribution, which are examined in this one.
First, the relation between the path length, $d_{i}$, and the number of dominant paths, $N$, is explored. To this end, the probability distribution of the path length conditioned on $N$, $P_{r}\left(d_{i}|N\right)$ is estimated. Obtained results are given in Fig. 6, which gives additional details on the already observed bimodal behavior of the EPDF of $d_{i}$. As seen, there are almost no paths with lengths in the range from $1000-3000$ m, approximately, except for the case where $N$ is very large, in which some paths with lengths in this range can be found. The values of $d_{i}$ are concentrated in the low length region and in the region around 3000 m, being (approximately) uniformly distributed from zero (or almost zero) in the low length region. Furthermore, the range of path lengths in the low region increases with the number of dominant paths.
Figure 6: Conditional probability of the path length, $d_{i}$, given the number of dominant paths, $N$, estimated over the set of measured channels.
Next, the relation between the absolute value of the path gain and the path length is assessed by estimating the conditional PDF of $|g_{i}|$ given $d_{i}$, $f_{|g_{i}|}\left(|g_{i}|\big{|}d_{i}\right)$. Results are displayed in Fig. 7. Interestingly, $|g_{i}|$ tends to a uniform distribution in the interval $(0,1)$ for very low values of $d_{i}$. As the path length increases, the range of values of $|g_{i}|$ decreases until it tends, again, to a uniform one in the interval $(0,1)$ for path lengths around $3000$ m. There is a clearly noticeable gap for path lengths between $1000-2500$ (approximately), where the few existing paths have very low gains.
Figure 7: Conditional PDF of $|g_{i}|$ given the path length, $d_{i}$, estimated over the set of measured channels.
Finally, the correlation between the path gains is explored. Since the decimation process described in Section III.III-A yields a different number of paths for each measured channel, the following procedure is employed to allow comparing results from different channels. In order to distinguish the MPM model parameters before the decimation process is applied from the ones resulting after it, let us denote the latter by using tilded symbols. Hence, the original set of path lengths and path gains are expressed as $d_{n}$ and $g_{n}$, respectively, with $10\leq n\leq N-$ 1, where $N=$2554, and their corresponding sets after the decimation process as $\tilde{d}_{i}$ and $\tilde{g}_{i}$, respectively, with $0\leq i\leq\tilde{N}$, where $\tilde{N}$ is the number of dominant paths after the decimation process. These magnitudes are related as
$g_{n}=\begin{cases}\tilde{g}_{i}&\text{if }\exists i\;|\;\tilde{d}_{i}=d_{n}\\ 0&\text{otherwise}\end{cases}.$ (16)
In order to explore the correlation among the path gains, the deterministic autocorrelation function (ACF) of $|g_{n}|$ for each measured channel, $R_{|g_{n}|}(n)=\frac{1}{N}\sum_{m=0}^{N-n-1}|g_{n+m}||g_{m}|$, for $0\leq n\leq N$, is computed and normalized. Fig. 8 depicts the its average value (over the set of measured channels) and the envelope of the ACF of all channels. As seen, the ACF decreases very rapidly, with values below 0.5 for $n\geq 8$ and below 0.3 for $n\geq 40$, and slightly regrows for large values of $n$ due to the behavior already shown in Fig. 7.
Figure 8: Average and envelope of the normalized deterministic ACF of $|g_{n}|$ computed over the set measured channels.
V Conclusion
The MPM is one of the most common top-down models adopted in the literature about PLC. This paper explores the parameterization of this model, which was originally proposed for low-voltage power distribution networks in the band up to 20 MHz, for the indoor broadband scenario in the band up to 80 MHz. Since the number of parameters that result from the extension of the MPM model to this context is extremely large, some of them have been imposed on the basis of common sense intuition, others, such as the attenuation parameters and the set of initial path lengths, have been chosen according to physical assumptions, and the remaining ones are analytically retrieved by means of a WLS fitting.
Since the number of parameters that result after the WLS fitting is still excessively large, only the dominant paths that actively contribute to synthesize the CFR have been selected according to a proposed decimation procedure that yields a more compact model. The implemented selection method determines the number of significant paths subject to a constraint on the maximum value of the NRMSE.
Once the MPM parameters have been computed according to the proposed strategy, they are related to the characteristics of the actual CFRs like the frequency selectivity and the average channel gain. Suitable distributions that fit the empirical PDF of each parameter are also given. Finally, the relation between the most relevant parameters of the model is explored. Provided results will aid the development improved channels models for indoor broadband PLC.

### v4

@@
-autocorrelation function CFR channel frequency response
+autocorrelation function AWGN additive white Gaussian noise
 Abstract
 This paper focuses on the parameterization of the multipath propagation model (MPM) for indoor broadband power line communications (PLC), which up to now has been established in an heuristic way.
-The MPM model was originally proposed for outdoor channels in the band up to 20 MHz, but its number of parameters becomes extremely large when used to model indoor channel frequency responses (CFR), which are much more frequency-selective than outdoor ones, and the band is extended to 80 MHz.
+The MPM model was initially proposed in the PLC context for outdoor channels in the band up to 20 MHz, but its number of parameters becomes extremely large when used to model indoor channel frequency responses (CFR), which are much more frequency-selective than outdoor ones, and the band is extended to 80 MHz.
 This work proposes a fitting procedure that addresses this problem.
 It allows determining the model parameters that yield the best fit to each channel of a large database of single-input single-output (SISO) experimental measurements acquired in typical home premises of different European countries.
-Then, the properties of the MPM parameters are analyzed.
+Then, the statistics of the MPM parameters are analyzed.
 The study unveils the relation between the model parameters and the main characteristics of the actual CFR like the frequency selectivity and the average attenuation.
 It also estimates the probability density function (PDF) of each parameter and proposes a fitting distribution for each of them.
-Moreover, the relationship among the main parameters of the model is also explored.
+Moreover, the relationship among the main parameters of the model, as well as their impact on the performance of PLC communication systems are also explored.
 Provided results can be helpful for the development of MPM-based models for indoor broadband PLC.
-Index Terms:
-Power line communications (PLC), channel model, multipath propagation model, channel frequency response, statistical characterization, probability density function.
 I Introduction
@@
 The power line communication (PLC) technology is becoming an established solution since it exploits the existing wired infrastructure to convey high-speed data content, thus saving costs and deployment time.
-Furthermore, the latest standards and devices include advanced signal processing and enhanced transmission techniques, such as the exploitation of all the deployed network conductors and the bandwidth extension, that improve even further the achievable data rate, guaranteeing transmission rates in the order of gigabit per second at the physical layer [1, 2, 3].
-This high performance level motivates a renewed interest in developing novel and updated PLC solutions, especially for home networking and smart grid applications [4, 5, 6, 7, 8].
+Furthermore, the latest standards and devices include advanced signal processing and enhanced transmission techniques, such as the exploitation of all the deployed network conductors and the bandwidth extension, that improve even further the achievable data rate, guaranteeing transmission rates in the order of gigabit per second at the physical layer [Yonge13, G.9960, IEEE_P1901].
+This high performance level motivates a renewed interest in developing novel and updated PLC solutions, especially for home networking and smart grid applications [Bush2013, Prasad2020, Righini2021, Canete2023, Cortes23bis].
 In order to support the growth of PLC technology, it is fundamental to provide effective channel models, which are able to numerically emulate the properties of actual PLC networks avoiding on field tests.
-Two basic approaches can be tackled in defining a model, i.e., bottom-up and top-down, and both of them can follow a deterministic or a statistical strategy [9, 10, 11, 12].
+Two basic approaches can be tackled in defining a model, i.e., bottom-up and top-down, and both of them can follow a deterministic or a statistical strategy [Zimmermann02, Tonello07bis, Canete2011, Povedano23].
 A deterministic strategy typically considers a specific network configuration, while a statistical strategy attempts to introduce some variability by means of statistical distributions.
@@
 Typically, it provides a faithful channel representation, although it turns out to be a computationally onerous procedure.
-Some examples concerning the single-input-single-output (SISO) channel modeling can be found in [13, 14], which propose a deterministic approach, and in [15, 16, 17, 18, 11], which follow a statistical approach.
-Recently, the models discussed in [19] and [20] extends the SISO bottom-up strategy to the multiple-input multiple-output (MIMO) case.
+Some examples concerning the single-input-single-output (SISO) channel modeling can be found in [Canete02, Galli06], which propose a deterministic approach, and in [Philipps99, Philipps00, Esmailian03, Tonello11, Canete2011], which follow a statistical approach.
+The models discussed in [Versolatto11] and [Corchado16] extend the SISO bottom-up strategy to the multiple-input multiple-output (MIMO) case.
 Conversely, the top-down approach is an empirical method that fits an analytic function to a measurements database.
 This strategy facilitates the statistical extension by adopting a certain distribution for the parameters that describe the theoretical function.
-One of the first and more common SISO models is the multipath propagation model (MPM) proposed in [9] and was later statistically extended in [21].
-Extensions of the SISO formulation to the MIMO are presented in [22] and [23].
-The PLC channel response also exhibits a short-term variation synchronous with the mains voltage, which allows modelling it as a linear periodically time-varying (LPTV) system [24].
-However, since the time variation is quite small compated to the duration of the impulse response, i.e., the channel response is underspread, the slow variation approximation can be made and the channel response can be modeled as a series of time-invariant responses that repeat with the period of the mains signal.
+The multipath propagation model (MPM) is an example of this approach, and its validity in both outdoor and indoor PLC scenarios is supported by the fact that power line cables are electrically long at the considered frequencies, resulting in a received signal that consists of multiples echoes caused by impedance mismatches at multiple points in the network.
+One of the first and more common MPM SISO models was proposed in [Zimmermann02] and later statistically extended in [Tonello12].
+Extensions of the SISO formulation to the MIMO case are presented in [Canova2010] and [Pagani16].
+The PLC channel response also exhibits a short-term variation synchronous with the mains voltage, which allows modelling it as a linear periodically time-varying (LPTV) system [Canete06].
+However, since the time variation is quite small compared to the duration of the impulse response, i.e., the channel response is underspread, the slow variation approximation can be made and the channel response can be modeled as a series of time-invariant responses that repeat with the period of the mains signal.
 Hence, the MPM model can be applied to each of these time-invariant responses.
 I-A Contribution
-The top-down MPM model formulated in [9] has been widely used in the PLC scenario [25, 10, 26, 27, 28].
-However, it was proposed for outdoor PLC channels in the band up to $20$ MHz.
-Its statistical extension to indoor channels made in [29] and [21] was based on heuristic approaches and no rigorous measurement-based analysis of this model for indoor scenarios in the band up to 80 MHz has been found in the literature.
+The MPM model has been widely used in PLC scenarios [Zimmermann02, OperaChannel05, Tonello07bis, Chien15, Huang2012, Gay2016].
+Its validity for outdoor channels in the frequency band up to $20$ MHz was empirically assessed in [Zimmermann02], which provided the model parameters that yielded the best fit to a set of measured channels.
+Statistical extensions to indoor channels were made in [Tonello07] and [Tonello12], which proposed channel generators that assumed certain a priori probability distributions for the model parameters based on heuristic approaches.
+However, a rigorous measurement-based analysis neither of the MPM capacity to model indoor scenarios in the band up to 80 MHz nor of the statistical distributions of the model parameters has been found in the literature.
 In this context, we make three main contributions:
   • We propose a method to determine the parameters of the MPM model for indoor broadband PLC channels in the band up to $80$ MHz.
-It addresses the problem of the extremely large number of parameters that results when the MPM model in [9] is extended to indoor scenarios and that are a consequence of the much more frequency selective behavior of indoor channel frequency responses and of the bandwidth extension to $80$ MHz.
-Since interference with commercial frequency modulation (FM) broadcasting is important above $87$ MHz in many regions.
+It addresses the problem of the extremely large number of parameters that results when the MPM model in [Zimmermann02] is extended to indoor scenarios and that are a consequence of the much more frequency selective behavior of indoor channel frequency responses and of the bandwidth extension to $80$ MHz.
   • The proposed method has been used to perform a channel-by-channel fitting, determining the set of MPM parameters that yields the best approximation to each of the channels of a large database consisting of 426 CFRs measured in indoor scenarios of different European countries.
   • We give an in-depth analysis of the resulting MPM parameters, showing their relation to the main channel characteristics like the frequency selectivity and the average attenuation, assessing their statistical behavior and the relation between them.
-Unlike the a posteriori parameters inference discussed in previous works, this study unveils the intrinsic properties of the MPM parameter and can be helpful for the development of novel MPM models or the refinement of existent ones.
+Unlike the a posteriori parameters inference discussed in previous works, this study unveils the intrinsic properties of the MPM parameter and can be helpful for the development of novel MPM models or the refinement of existing ones.
+A discussion in this regard, and a study of the relationship between the main MPM parameters and the performance of indoor PLC systems are also provided.
 I-B Organization
@@
 The properties of the computed parameters, such as their relation to the channel characteristics, statistical behavior and relation between them are discussed in Section IV.
-Finally, conclusions follow in Section V.
+The application of the proposed results to a further development of MPM-based channel generators and the analysis of the influence of the model parameters in the performance of PLC systems is discussed in Section V.
+Finally, conclusions follow in Section VI.
 II Background: The Multipath Propagation Model
 The MPM follows a top-down approach and describes the multipath propagation of the power signal into the PLC medium that is due to the line discontinuities, such as branches, or unmatched loads.
-The analytic formulation of the CFR can be expressed as [9]
+The analytic formulation of the CFR can be expressed as [Zimmermann02]
 $H\left(f\right)=A\sum_{i=0}^{N-1}{g_{i}}e^{-(a_{0}+a_{1}f^{K})d_{i}}e^{-j2\pi fd_{i}/v},$ (1)
-where $a_{0}$ and $a_{1}$ are the attenuation coefficients, $g_{i}$ and $d_{i}$ are the path gain and the path length of the $i$-th path [^1] For the sake of notation simplicity in the subsequent fitting process, paths have been indexed as $10\leq i\leq N-$ 1, instead of $1\leq i\leq N$ as done in [9]., respectively.
-The coefficient $g_{i}$ can be interpreted as the product of transmission and reflection coefficients by restricting $|g_{i}|\leq 1$ and by introducing the normalization coefficient $A>0$, which allows adjusting the attenuation of the resulting CFR [9].
+where $a_{0}$ and $a_{1}$ are the attenuation coefficients, $g_{i}$ and $d_{i}$ are the path gain and the path length of the $i$-th path [^1] For the sake of notation simplicity in the subsequent fitting process, paths have been indexed as $0\leq i\leq N-1$, instead of $1\leq i\leq N$ as done in [Zimmermann02]., respectively.
+The coefficient $g_{i}$ can be interpreted as the product of transmission and reflection coefficients by restricting $|g_{i}|\leq 1$ and by introducing the normalization coefficient $A>0$, which allows adjusting the attenuation of the resulting CFR [Zimmermann02].
 The term $N$ denotes the number of paths and $K$ is the exponent of the attenuation factor related to the cable features.
 This work uses $N=2554$ (its rationale will be justified in Section III.III-B), which is much larger than the values used in outdoor channels.
-Typical values of $K$ range between 0.5 and 1 [9], hence $K=1$ is assumed for the rest of this paper.
-The parameter $\nu=c\epsilon_{r}$ is the propagation speed of light in the cables, where $c$ is the speed of light in the vacuum and $\epsilon_{r}$ is the relative dielectric constant of the insulator which envelops the conductors, which has been set to $\epsilon_{r}=1.5$ in order to take into account the non-uniformity of the dielectric [9].
-III Determination of the MPM Parameters
+Typical values of $K$ range between 0.5 and 1 [Zimmermann02], hence $K=1$ is assumed for the rest of this paper.
+The parameter $\nu=c\epsilon_{r}$ is the propagation speed of light in the cables, where $c$ is the speed of light in the vacuum and $\epsilon_{r}$ is the relative dielectric constant of the insulator which envelops the conductors, which has been set to $\epsilon_{r}=1.5$ in order to take into account the non-uniformity of the dielectric [Zimmermann02].
+Expression (1) is a truncated version of the actually infinite number of forward- and backward-traveling waves caused by the reflections due to impedance mismatch.
+The rationale for this truncation is that the amplitude of these waveforms decreases at each reflection.
+Hence, those that have experienced a larger number of reflections would have negligible amplitude and can be ignored.
+It must be emphasized that the path lengths, $d_{i}$, correspond to the actual distances traveled by the referred waves.
+For instance, in a simple network consisting of a voltage source and a mismatched transmission line of length $\mathcal{L}$ terminated in a load impedance, the distances travelled by the received waves are $d_{i}=(2i+1)\mathcal{L}$, with $i=0,1,2,\ldots$ [Paul94, Ch.
+5].
+III Method to Determine the MPM Parameters
 In order to provide a statistical characterization of the parameters that describe the MPM formulation in (1), the first step is to compute the MPM parameters that give the best fitting of this analytic expression to each of the channels of the measurement database.
-The latter consists of $426$ indoor CFRs measured in different European countries using a vector network analyzer.
-CFRs have $M=1262$ frequency samples given by $f_{m}=f_{0}+m\Delta f$, with $f_{0}=1.0$ MHz, $10\leq m\leq M-$ 1 and $\Delta f=62.5978$ kHz.
-Part of this collection was previously used in [23] and [12], and the results are consistent with the ones obtained from other extensive measurements in Italy that were characterized in [30].
-Since the model is described by many parameters, the following assumptions must be adopted in order to simplify the fitting strategy.
-First, the attenuation coefficients $a_{0}$ and $a_{1}$ are estimated through the robust regression fit of each channel measurement.
-This linear attenuation profile can assumed to be essentially determined by the longest path of the MPM.
-Hence, we consider the model in (1) with a single path of the largest length, $L$ (its value will be given in Section III.III-B).
-Since $A$ is a normalization factor used to force $|g_{i}|\leq 1$, when a single path is considered it happens that $|A\cdot g_{i}|=1$.
-Accordingly, the robust regression fit of the longest path in (1) can be expressed as
+It consists in a set of $426$ indoor CFRs corresponding to differential transmission and reception between the phase and neutral conductors (P-N), out of which $205$ have been acquired in Spain and $221$ where measured in the ETSI STF 410 MIMO PLC campaign described in [Pagani16].
+Measured CFRs consists of $M=1262$ frequency samples given by $f_{m}=f_{0}+m\Delta f$, with $f_{0}=1.0$ MHz, $0\leq m\leq M-1$ and $\Delta f=62.5978$ kHz.
+The characteristics of the measured channels are similar to those of the extensive campaign carried out in Italy [Tonello14] and to the ones measured in the United States and reported in [Galli11].
+They also exhibit well-known properties of indoor PLC channels, such as the lognormal distribution of the delay spread, which has a mathematical foundation [Galli11].
+The features of the employed channels are also coherent with the ones in [Pagani16].
+However, it is worth mentioning that while the $351$ channels used in the latter to derive a MIMO channel model were also acquired in the referred ETSI campaign, they correspond to the P, N, E and CM (common mode) receiver ports described in [ETSI_TR_101_562-3], while in this work the conventional SISO differential P-N links are employed.
+Since the model is described by many parameters, the following procedure is used to simplify the fitting strategy.
+First, the attenuation coefficients $a_{0}$ and $a_{1}$ are estimated.
+To this end, we consider an MPM consisting on a single path.
+By setting $N=1$ in (1), it happens that $|A\cdot g_{i}|=1$, since $A$ is a normalization factor used to force $|g_{i}|\leq 1$.
+Furthermore, $N=1$ yields a linear attenuation profile from which the attenuation coefficients $a_{0}$ and $a_{1}$ can be estimated through the regression fit of each channel measurement.
+This linear attenuation profile can be assumed essentially determined by the longest path of the MPM, which is denoted by $d_{1}=L$ (its value will be given in Section III-B).
+The regression fit can be expressed as
 $10\log_{10}|H(f)|^{2}=-2(a_{0}+a_{1}f)L\cdot 10\log_{10}e=\alpha_{0}+\alpha_{1}f,$ (2)
-where $\alpha_{0}$ and $\alpha_{1}$ are the coefficients of the robust regression fit. Thus, the attenuation coefficients are computed as
+where $\alpha_{0}$ and $\alpha_{1}$ are the coefficients of the regression fit. Thus, the attenuation coefficients are computed as
 $a_{0}=-\frac{\alpha_{0}}{20L\log_{10}e},\quad a_{1}=-\frac{\alpha_{1}}{20L\log_{10}e}.$ (3)
+The presented procedure does not impose a rigid constraint on $a_{0}$, as there are multiple values of $a_{0}$, $g_{i}$ and $d_{i}$ that yield the same $g_{i}e^{-a_{0}d_{i}}$.
+Hence, if the longest path of the considered channel is shorter than $L$, the small value of $a_{0}$ resulting from the employed procedure can be compensated with larger values of $g_{i}$.
+Regarding $a_{1}$, which is related to the slope of the attenuation profile, its value is generally small, since the attenuation in indoor PLC channels does not exhibit the remarkable frequency dependent profile of outdoor channels [Cortes23, Fig.5], as involved distances are much shorter and attenuation is mainly due to multipath propagation and not to the skin effect.
+The validity of these assumptions will be corroborated later by the fitting results.
 Second, the path lengths are discretized yielding uniformly spaced distances,
 $d_{i}=i\frac{L}{N}\quad\text{with }i=0,...,N-1.$ (4)
-This is a viable assumption since indoor PLC networks typically exhibit a complex topology with a great number of branches.
-This results into approximately uniformly distributed distances (spaced $1.25$ m for the values of $L$ and $N$ considered in this work) traveled by the signal and its reflections from the transmitter side towards the receiver outlet.
-Afterwards, the path gains, $g_{i}$ that give the best fitting to each each measured channel have to be determined.
+It should be noted that actual path lengths are inherently discrete, as the lengths of the cable sections through which the waves propagate (forward and backward) are discrete quantities.
+Since indoor PLC networks have complex topologies with numerous branches whose lengths are unknown a priori, the uniform distribution appears to be the most appropriate choice.
+This assumption defines a grid of potential path lengths.
+Actual path lengths would be rounded to these grid elements.
+For the values of $L$ and $N$ used in this work (which will be given in Section III-B), the spacing is $1.25$ m.
+This means that the maximum rounding error of a path length would be $1.25/2=0.625$ m, which seems to be adequate even for small power line networks.
+Furthermore, it must be emphasized that the uniform grid is used as a starting point, those $d_{i}$ that do not correspond to actual (rounded) propagation distances within the considered network will have $g_{i}\approx 0$ and will be pruned in subsequent steps of the procedure.
+Afterwards, the path gains, $g_{i}$, that give the best fitting to each measured channel have to be determined.
 This could be done by means of a LS estimation that minimizes the root mean square error (RMSE) w.r.t.
@@
 $\text{NRMSE}=\sqrt{\frac{1}{M}\sum_{m=0}^{M-1}{\frac{|H(f_{m})-\hat{H}(f_{m})|^{2}}{|H(f_{m})|^{2}}}}.$ (6)
-The $g_{i}$ can be obtained by solving a weighted (WLS) problem formulated by expressing (1) in matrix form as follows,
+The gains $g_{i}$ can be obtained by solving a weighted (WLS) problem formulated by expressing (1) in matrix form as follows,
 $\underbrace{\begin{bmatrix}H(f_{0})\\ \vdots\\ H(f_{M-1})\\ H^{*}(f_{M-1})\\ \vdots\\ H^{*}(f_{0})\\ \end{bmatrix}}_{\mathbf{h}}=\underbrace{\begin{bmatrix}P_{0,0}&\dots&P_{0,N-1}\\ \vdots&&\vdots\\ P_{M-1,0}&\dots&P_{M-1,N-1}\\ P^{*}_{M-1,0}&\dots&P^{*}_{M-1,N-1}\\ \vdots&&\vdots\\ P^{*}_{0,0}&\dots&P^{*}_{0,N-1}\\ \end{bmatrix}}_{\mathbf{P}}\underbrace{\begin{bmatrix}g_{0}\\ \vdots\\ g_{N-1}\end{bmatrix}}_{\mathbf{g}}$ (7)
@@
 $j=\operatorname*{arg\,min}_{i}{\left\{\sum_{m}|g_{i}|e^{-(a_{0}+a_{1}f_{m})d_{i}}\right\}}.$ (10)
-The rationale for this criterion is that the actual contribution of a path to the overall CFR is essentially determined by the factor $|g_{i}|\cdot\exp(-(a_{0}+a_{1}f)d_{i}$ in the MPM in (1).
-The decimation procedure continues while the NRMSE (dB) remains lower than a given threshold, which has been set to $-20$ dB.
+The rationale for this criterion is that the contribution of a path to the overall CFR is determined by $|g_{i}|e^{-(a_{0}+a_{1}f)d_{i}}$ and not solely by the path length.
+Indeed, results presented in Section IV-B corroborate that the longest path is not generally the one with the smallest contribution.
+The decimation procedure continues while the NRMSE (dB) remains lower than a given threshold, which has been set to 20-20 dB.
 This ensures that the channel computed with a reduced set of paths still represents a very good approximation to the corresponding measured channel.
@@
 To this end, the problem in (7) is reformulated by removing the $j$-th column of the matrix $\mathbf{P}$ and the new path gains are obtained.
-Table I summarizes the presented fitting and decimation procedure applied to each measured channel, where the expressions to obtain the values of $L$ and $N$ will be given in Section III.III-B.
+Table I summarizes the presented fitting and decimation procedure applied to each measured channel, where the expressions to obtain the values of $L$ and $N$ will be given in Section III-B.
 When the decimation procedure ends, the path gains obtained for the considered measured channel are normalized by their maximum absolute value.
@@
 6: Compute the NRMSE using (6)
-7: while NRMSE (dB)$<\text{NRMSE}_{\text{threshold}}$ do
+[…]
    8: Obtain $j$ using (10)
@@
    12: Compute the NRMSE using (6)
-     13: Update $N=N-1$
+   13: Update N=N1N=N-1
 14: end while
@@
 17: Update $\mathbf{\widehat{h}}=\mathbf{P}\mathbf{g}$
-The number of initial paths considered for the model at the pre-decimation point is $N=2554$ (as justified in the next subsection).
-This gives average and standard deviation values of the NRMSE (computed over the set of measured channels) equal to $\overline{\textrm{NRMSE}}=-36.79$ dB and $\sigma_{\textrm{NRMSE}}=7.99$ dB, respectively.
-After the application of the proposed decimation procedure, the number of paths is reduced to an average value (over the whole set of fitted channels) of 217.84, i.e., less than 10% of the initial value, and the corresponding average and standard deviation values of the NRMSE (dB) are $\overline{\textrm{NRMSE}}=-20.05$ dB and $\sigma_{\textrm{NRMSE}}=0.0678$ dB, respectively.
-Hence, the proposed decimation process achieves a good trade-off between modeling complexity and accuracy.
-For illustrative purposes, Fig.1 shows the amplitude and unwrapped phase of a measured CFR, $H(f)$, and its corresponding fitted response, $\hat{H}(f)$, obtained with $N=183$ dominant paths determined using the proposed decimation process.
+Table II summarizes the results obtained when the aforementioned decimation path procedure is applied to fit the measured channels in the frequency bands up to $20$ MHz and up to $80$ MHz.
+The former is included for comparison with the seminal work in [Zimmermann02], which applied the MPM to outdoor scenarios in the frequency band up to $20$ MHz.
+The number of initial paths at the pre-decimation stage is $N=640$ for the $20$ MHz bandwidth and $N=2554$ for the $80$ MHz (as justified in the next subsection).
+As expected, the average number of paths after the decimation process is much smaller in the $20$ MHz case (less than half) than in the $80$ MHz one.
+Also, its variance $\sigma_{N}$ is reduced.
+However, the number of paths is still much larger than in outdoor scenarios [Zimmermann02], where channels are less frequency selective because of the reduced number of derivations and increased attenuation (due to longer distances) that reduce the destructive effect of the multipath propagation.
+Regarding the efficiency of the decimation procedure, the number of paths is reduced to an average value (over the whole set of fitted channels) of $107.97$ in the $20$ MHz case and to $217.84$ in the $80$ MHz bandwidth.
+Interestingly, the former is about 16.816.8% of the initial $N$, while the latter is lower than 1010% of the initial $N$.
+The higher efficiency in the $80$ MHz band is a consequence of the lower frequency selectivity of indoor PLC channels above 30 MHz [Cortes23bis].
+[…]
+The running time of the iterations of the decimation procedure [^3] Executed by MATLAB R2024b in a computer with Intel core i9-10900 CPU at 2.80 GHz and 32 GB of RAM.
+quadratically decreases from approximately $7$ s for the first one to less than $0.1$ s for iterations above $2000$, since the size of the matrices in (9) get smaller as the number of decimated paths increases.
+Nevertheless, it must be emphasized that the presented decimation process is not subject to any real-time constraint, as it might be the case of a real-time MPM channel generator.
+Table II: Summary of the dominant path selection procedure for channels up to $20$ MHz and $80$ MHz
+Frequency band
+Magnitude 20 MHz 80 MHz
+Value of $N$ before decimation 640 2554
+$\overline{\textrm{NRMSE}}$ (dB) before decimation -30.77 -36.79
+$\sigma_{\textrm{NRMSE}}$ (dB) before decimation 7.69 7.99
+Average $N$ after decimation 107.97 217.84
+$\sigma_{N}$ (dB) after decimation 55.50 134.58
+$\overline{\textrm{NRMSE}}$ (dB) after decimation -20.11 -20.05
+$\sigma_{\textrm{NRMSE}}$ (dB) after decimation 0.1537 0.0687
+[…]
+Presented results shows that the proposed decimation process achieves a good trade-off between modeling complexity and accuracy.
+For illustrative purposes, Fig.
+2 shows the amplitude and unwrapped phase of a measured CFR, $H(f)$, and its corresponding fitted response, $\hat{H}(f)$, obtained with $N=183$ dominant paths determined using the proposed decimation process.
 As can be observed, the measured trace is almost veiled by the estimated one, which highlights the notable accuracy of the fitting process.
@@
 III-B Relation Among L, M, and N
-The maximum length $L$, the maximum number of paths $N$, and the number of frequency samples $M$ depend among each other and must satisfy certain constraints.
+The maximum length $L$, the maximum number of paths $N$ depend on the considered frequency band and the frequency resolution of the measured CFR.
+Hence, they are also related to the number of frequency samples, $M$, of the latter and must satisfy certain constraints which are now derived.
 Since only a discrete and finite number of frequencies is available, the expression in (1) given for the continuous frequency domain can be rewritten for the discrete domain as
@@
 Therefore, the relation between $L$ an the frequency samples $M$ is given by
-$\Delta f=\frac{f_{M-1}-f_{0}}{M-1}=\frac{1}{N\frac{L}{N\nu}}=\frac{\nu}{L}\Rightarrow L=\frac{(M-1)\nu}{f_{M-1}-f_{0}}$ (12)
-Since $f_{M-1}=79.9358$ MHz and $f_{0}=1$ MHz, $M=1262$ and the propagation speed is $\nu=2\cdot 10^{8}$m/s, the maximum possible length that can be assumed is $L=3195$ m.
-The limit of more than 3 kilometers turns out to be a maximum length that is long enough to describe any generic in-home network topology.
-Moreover, always referring to the DFT representation of the CFR reported in (11), the DFT frequency period, i.e., $N\Delta f$, must be greater or equal than two times the maximum channel frequency, i.e., $f_{M-1}$, in order to avoid aliasing problems according to the sampling theorem.
+$\Delta f=\frac{f_{M-1}-f_{0}}{M-1}=\frac{1}{N\frac{L}{N\nu}}\Rightarrow L=\frac{\nu}{\frac{f_{M-1}-f_{0}}{M-1}}=\frac{\nu}{\Delta f}.$ (12)
+Expression (12) can be also derived from physical considerations.
+The time it takes for an input impulse to the channel to reach the receiver following the longest path (i.e., the one with the largest number of reflections) is $\tau=L/\nu$.
+The maximum duration of the channel impulse response that can be obtained from a sampled version of the CFR at multiples of $\Delta f$ while avoiding time-domain aliasing is $\tau=1/\Delta f$.
+Equating both expressions yields $L=\nu/\Delta f$.
+For $f_{M-1}=79.9358$ MHz and $f_{0}=1$ MHz, $M=1262$ and the propagation speed is $\nu=2\cdot 10^{8}$m/s, the maximum possible length that can be assumed is $L=3195$ m.
+Since physical distances in indoor PLC networks are on the order of tens or few hundreds of meters, the limit of over $3$ kilometers is large enough to account for the longest distance traveled by all forward- and backward-propagating waves received with significant power.
+Referring to the DFT representation of the CFR reported in (11), the DFT frequency period, i.e., $N\Delta f$, must be greater or equal than two times the maximum channel frequency, i.e., $f_{M-1}$, in order to avoid aliasing.
 Thus, the relationship among $N$ and $L$ turns out to be
-$N\Delta f=\frac{1}{\tau}\geq 2f_{M-1}\Rightarrow N\geq\frac{2f_{M-1}L}{\nu}.$ (13)
+$N\Delta f=\frac{1}{\tau}\geq 2f_{M-1}\Rightarrow N\geq\frac{2f_{M-1}L}{\nu}=\frac{f_{M-1}}{\Delta f}.$ (13)
+As seen, $N$ is determined by the considered bandwidth and the frequency resolution of the measurements.
 If the aforesaid values for $f_{M-1}$ and $\nu$ are considered, the relation in (13) becomes $N\geq 0.8L$.
@@
 This provides the last relationship, expressed by
-$d_{1}=\frac{L}{N}\geq\frac{\lambda_{min}}{2}\Rightarrow N\leq\frac{2f_{M-1}L}{\nu}$ (15)
-It can be noted that, given the above mentioned specifications, as well as the relations (13) and (15), the relationship among $L$ and $N$ must hold with equality, i.e., $N=0.8L=2554$ for the considered database.
-IV Analysis of the MPM Parameters
+$d_{1}=\frac{L}{N}\geq\frac{\lambda_{min}}{2}\Rightarrow N\leq\frac{2f_{M-1}L}{\nu}=\frac{f_{M-1}}{\Delta f}.$ (15)
+Expression (15) can be also derived from physical considerations.
+Assuming uniform distances, the time difference between the closest received paths is $\frac{L/N}{\nu}$.
+In order for these paths to be resolvable, this time must be greater or equal than the minimum sampling period, $T_{s}=\frac{1}{2f_{M-1}}$.
+This yields $\frac{L/N}{\nu}\geq\frac{1}{2f_{M-1}}$, which equals (15).
+Given the specifications of the measured CFR, as well as the relations (13) and (15), the relationship among $L$ and $N$ must hold with equality, i.e., $N=0.8L=2554$ for the considered database.
+IV Statistical Analysis of the MPM Parameters
 This section analyzes the statistical properties of the MPM parameters obtained according to the fitting procedure detailed in Section III for the overall set of measured channels.
-First, the relation of some model parameters to the main magnitudes used to characterize the channel response, such as the delay spread and the average channel gain.
+First, the relation of some model parameters to the main quantities used to characterize the channel response, such as the delay spread and the average channel gain.
 Then, the PDF of the MPM parameters is estimated and, finally, the statistical relation between the model parameters is studied.
 IV-A Relation to the Channel Characteristics
-The parameters $a_{0}$ and $a_{1}$ where obtained from the robust regression fit of the amplitude response of the measured channels, with $a_{1}$ being related to the slope of the attenuation profile and $a_{0}$ with the y-intercept.
+The parameters $a_{0}$ and $a_{1}$ were obtained from the regression fit of the amplitude response of the measured channels, with $a_{1}$ being related to the slope of the attenuation profile and $a_{0}$ with the y-intercept.
 Hence, the latter parameter has to be somehow related to the average channel gain, defined as $G\textrm{ (dB)}=\frac{1}{M}\sum_{m=0}^{M-1}20\log_{10}{|H(f_{m})|}$}.
 This relation can be clearly observed in Fig.
-2 (a), where the scatter plot of $a_{0}$ vs $G$ (dB) is depicted, along with its robust fit regression line, whose expression is given in Table II.
+3 (a), where the scatter plot of $a_{0}$ vs $G\;\textrm{(dB)}$} is depicted, along with its robust fit regression line, whose expression is given in Table III.
 As expected, channels with larger average gains need lower values of $a_{0}$ to be synthesized.
-Figure 2: Scatter plot of (a) the value of $\bm{a_{0}}$ vs the average channel gain of each measured channel, $\bm{G}$ (dB); (b) the number of dominant paths vs the average gain and the delay spread of the channels; (c) value of $\bm{A}$ vs the average channel gain of each measured channel.
+Figure 3: Scatter plot of (a) the value of $a_{0}$ vs the average channel gain of each measured channel, $G$ (dB); (b) the number of dominant paths vs the average gain and the delay spread of the channels; (c) value of $A$ vs the average channel gain of each measured channel.
 The number of dominant paths needed to fit a given CFR is related to the frequency selectivity of the channel, e.g., a single path suffices to synthesize a flat channel.
-The frequency selectivity of a channel is usually quantified by means of the coherence bandwidth or its inversely related magnitude, the delay spread (also referred to as multipath spread) [31].
-In PLC these magnitudes are known to be related to the average channel gain [8]: channels with lower average channel gains tend to have larger delay spreads.
+The frequency selectivity of a channel is usually quantified by means of the coherence bandwidth or its inversely related magnitude, the delay spread (also referred to as multipath spread) [Molish2011].
+In PLC these magnitudes are known to be related to the average channel gain [Cortes23bis]: channels with lower average channel gains tend to have larger delay spreads.
 Hence, these magnitudes are expected to be related to the number of dominant paths, $N$.
@@
 3 (b) illustrates this relation by showing the scatter plot of the number of dominant paths vs the average channel gain and the delay spread of the corresponding measured channel.
-Its robust regression plane, whose expression is given in Table II, is also depicted for reference.
-The goodness of this fit is confirmed by the NRMSE between the actual and predicted values of $N$, which is $-22$ dB.
+Its robust regression plane, whose expression is given in Table III, is also depicted for reference.
+The goodness of this fit is confirmed by the NRMSE between the actual and predicted values of $N$, which is 22-22 dB.
 The value of the normalization parameter $A$ is also related to the average channel gain, as shown in Fig.
 3 (c), which depicts the scatter plot of $A$ vs $G$ (dB) and its corresponding exponential fitting function.
-The parameters of the latter, which have been obtained by means of a LS fitting, are given in Table II.
+The parameters of the latter, which have been obtained by means of a LS fitting, are given in Table III.
 As observed, the lower the average channel gain, the lower the value of $A$ needed to synthesize the channel.
-Table II: Parameters of the fitting expressions displayed in Fig. 2.
+Table III: Parameters of the fitting expressions displayed in Fig. 3.
 Fitting expression $\bm{\alpha}$ $\bm{\beta}$ $\bm{\gamma}$
-[…]
-[…]
-$\widehat{A}=\alpha\cdot e^{\beta\cdot G\textrm{ (dB)}}$ $8.2517\cdot 10^{-1}$ $1.0636\cdot 10^{-1}$
+$\widehat{a}_{0}=\alpha+\beta\cdot G\textrm{ (dB)}$ $1.8669\cdot 10^{-4}$ $3.4066\cdot 10^{-5} -$
+$\widehat{N}=\alpha+\beta\cdot ds\textrm{ ($\mu$s)}+\gamma\cdot G\textrm{ (dB)}$ $40.4009$ $185.7535$ $4.5097$
+$\widehat{A}=\alpha\cdot e^{\beta\cdot G\textrm{ (dB)}}$ $8.2517\cdot 10^{-1}$ $1.0636\cdot 10^{-1}$ -
 IV-B Statistical Characterization
-This subsection estimates the chief moments and the PDF of the MPM parameters, computed for the overall set of measured channels, that have been obtained using the fitting procedure presented in Section IV.
+This subsection estimates the main moments and the PDF of the MPM parameters, computed for the overall set of measured channels, that have been obtained using the fitting procedure presented in Section III.
 For each empirical (estimated) PDF, the probability distribution that gives the best fit is given.
-To this end, the most common distribution functions are considered: beta, Birnbaum-Saunders, exponential, gamma, generalized extreme value (GEV), Gumbel, inverse Gaussian, logistic, log-logistic, log-normal, Nakagami, normal, Poisson, Rayleigh, Rician, $t$ location-scale and Weibull.
-The parameters of each distribution that yield the best fit to the set of MPM parameters are obtained by means of the maximum likelihood (ML) estimation assuming independent samples [32].
+To this end, the most common distribution functions are considered: beta, Birnbaum-Saunders, exponential, gamma, generalized extreme value (GEV), Gumbel, inverse Gaussian, logistic, log-logistic, lognormal, Nakagami, normal, Poisson, Rayleigh, Rician, $t$ location-scale and Weibull.
+For each distribution, the parameters that yield the best fit to the set of MPM parameters are obtained by means of the maximum likelihood (ML) estimation assuming independent samples [Bickel2015].
 Afterwards, the Anderson-Darling goodness-of-fit test is applied and the distribution yielding the highest $p$-value is selected.
@@
 For some parameters, more elaborate models consisting of mixture distributions have been considered.
-The mean and standard deviation of the MPM parameters are given in Table III.
-The fitting statistical distributions with their corresponding parameter values are summarized in Table IV.
+The mean and standard deviation of the MPM parameters are given in Table IV.
+The fitting statistical distributions with their corresponding parameter values are summarized in Table V.
 These reported results allow future replicability of this study and simplify the development of improved statistical models.
-Table III: Mean and standard deviation of the MPM parameters
+Table IV: Mean and standard deviation of the MPM parameters
 MPM parameter mean standard deviation
@@
 Path lengths ($d_{i}$) 343.9491 633.4392
-Figure 3: Empirical and fitted PDFs of the attenuation coefficients $\bm{a_{0}}$ and $\bm{a_{1}}$ and the normalization coefficient $\bm{A}$.
-Figure 4: Empirical and fitted PDFs of the number of dominant paths, $\bm{N}$, the modulus of the path gains $\bm{|g_{i}|}$ and the path lengths $\bm{d_{i}}$.
+Figure 4: Empirical and fitted PDFs of the attenuation coefficients $a_{0}$ and $a_{1}$ and the normalization coefficient $A$.
+Figure 5: Empirical and fitted PDFs of the number of dominant paths, $N$, the modulus of the path gains $|g_{i}|$ and the path lengths $d_{i}$.
 The empirical s of the MPM parameters are reported in Fig.
 4 and Fig.
-4.
+5.
 Regarding $a_{0}$ and $a_{1}$, whose EPDFs are show in Fig.
-3 (a) and (b), respectively, the distribution that gives the best fit is the GEV ($p$-value$=0.286$ for $a_{0}$ and 0.8941 for $a_{1}$), although the normal distribution also gives a high $p$-value (0.1967 for $a_{0}$ and 0.5664 for $a_{1}$).
-In practice, the GEV is a family of three types of extreme value distributions, namely type I, or Gumbel (or log-Weibull) [33], type II, or Fréchet, and type III, or Weibull distribution.
+4 (a) and (b), respectively, the distribution that gives the best fit is the GEV ($p$-value=0.286=0.286 for $a_{0}$ and 0.8941 for $a_{1}$), although the normal distribution also gives a high $p$-value (0.1967 for $a_{0}$ and 0.5664 for $a_{1}$).
+In practice, the GEV is a family of three types of extreme value distributions, namely type I, or Gumbel (or log-Weibull) [Gumbel1954], type II, or Fréchet, and type III, or Weibull distribution.
 This result makes sense since the GEV is the limit distribution of properly normalized maxima of a sequence of independent and identically distributed random variables.
@@
 Regarding the normalization coefficient, $A$, whose EPDF is displayed in Fig.
-3 (c), the Log-logistic gives the largest $p$-value (0.909) although, although the GEV and the log-normal also give high values ($p$-values $0.788$ and $0.4536$, respectively).
-The latter is not surprising, as $A$ is determined as the maximum value of the path gains, $|g_{i}|$, of each fitted channel.
+4 (c), the Log-logistic gives the largest $p$-value (0.909) although, although the GEV and the lognormal also give high values ($p$-values $0.788$ and $0.4536$, respectively).
+The latter is not surprising, as $A$ is determined as the maximum of the absolute value of the path gains, $|g_{i}|$, of each fitted channel.
 Regarding the probability distribution of the number of dominant paths, $N$, while this magnitude can only take non-negative integer values, discrete distributions like the binomial and Poisson ones give very poor fitting to the actual set of values.
@@
 This same approach has been also applied to fit the path length, $d_{i}$.
-By doing so, the GEV gives the best fit ($p$-value$=0.86752$) to the number of dominant paths, $N$, whose EPDF is displayed in Fig.
+By doing so, the GEV gives the best fit ($p$-value=0.86752=0.86752) to the number of dominant paths, $N$, whose EPDF is displayed in Fig.
 5 (a).
@@
 Fig.
-4 (b) shows that the continuous term of the EPDF can be fitted by a log-normal RV.
-Accordingly, the overall PDF can be expressed as $f(|g_{i}|)=\pi_{0}q(|g_{i}|\big{|}\mu,\sigma)+\pi_{1}\delta(|g_{i}|-1)$, where $q(\cdot\big{|}\mu,\sigma)$ denotes the PDF of a log-normal RV with parameters $\mu$ and $\sigma$ and $\delta(\cdot)$ is the Dirac delta function.
-The value of these parameters are given in Table IV.
+5 (b) shows that the continuous term of the EPDF can be fitted by a lognormal RV.
+Accordingly, the overall PDF can be expressed as $f(|g_{i}|)=\pi_{0}q(|g_{i}|;\mu,\sigma)+\pi_{1}\delta(|g_{i}|-1)$, where $q(\cdot;\mu,\sigma)$ denotes the PDF of a lognormal RV with parameters $\mu$ and $\sigma$ and $\delta(\cdot)$ is the Dirac delta function.
+The value of these parameters are given in Table V.
 The EPDF of $d_{i}$, displayed in Fig.
@@
 It has been found that the Weibull distribution gives a good fitting to the values of $d_{i}\leq 1500$ while the GEV, shifted and reflected about $d_{N-1}=\frac{N-1}{N}L$, adequately fits the values of $d_{i}>1500$.
-Hence, the PDF of $d_{i}$ can be modeled as $f(d_{i})=\pi_{0}g_{0}(d_{i}\big{|}\lambda_{0},k_{0})+\pi_{1}g_{1}(d_{N-1}-d_{i}\big{|}k_{1},\sigma_{1},\mu_{1})$, where $g_{0}(\cdot\big{|}\lambda_{0},k_{0})$ denotes the PDF of a Weibull distribution with scale and shape parameters $\lambda_{0}$ and $k_{0}$, respectively, and $g_{1}(\cdot\big{|}k_{1},\sigma_{1},\mu_{1})$ is the PDF of a GEV distribution with location, scale and shape parameters $k_{1}$, $\sigma_{1}$ and $\mu_{1}$, respectively, whose values are given in Table IV.
+Hence, the PDF of $d_{i}$ can be modeled as $f(d_{i})=\pi_{0}g_{0}(d_{i};\lambda_{0},k_{0})+\pi_{1}g_{1}(d_{N-1}-d_{i};k_{1},\sigma_{1},\mu_{1})$, where $g_{0}(\cdot;\lambda_{0},k_{0})$ denotes the PDF of a Weibull distribution with scale and shape parameters $\lambda_{0}$ and $k_{0}$, respectively, and $g_{1}(\cdot;k_{1},\sigma_{1},\mu_{1})$ is the PDF of a GEV distribution with location, scale and shape parameters $k_{1}$, $\sigma_{1}$ and $\mu_{1}$, respectively, whose values are given in Table V.
 As shown in Fig.
 5 (c), the PDF of the proposed mixture gives a good fitting to the empirical one.
-Table IV: Fitting distributions of the MPM parameters
+Figure 6: (a) Dominant path length vs dominant path index for some illustrative measured channels; (b) scatter plot of the median value of the dominant path length vs the delay spread for the set of measured channels; (c) scatter plot of the median value of the dominant path length vs the number of dominant paths for the set of measured channels.
+Table V: Fitting distributions of the MPM parameters
 MPM
@@
 GEV
-$k=-0.2744$
+k=0.2744k=-0.2744
 $\sigma=4.9492\cdot 10^{-4}$
@@
 GEV
-$k=-0.1781$
+k=0.1781k=-0.1781
 $\sigma=3.8980\cdot 10^{-12}$
@@
 Norm. coeff.
-$A$
+AA
 Log-logistic
@@
 dominant paths
-$N$
+NN
 GEV
@@
 4, there are no paths with intermediate distances.
-This behavior is related to the frequency selectivity of the channel: the larger the frequency selectivity, the large the distances of the paths required to synthesize it.
+This behavior is related to the frequency selectivity of the channel: the larger the frequency selectivity, the larger the distances of the paths required to synthesize is.
 This can be observed in Fig.
@@
 As observed, channels with higher delay spread values (larger frequency selectivity) need larger paths to be synthesized.
-Futhermore, it is also expected that the higher the number of dominant paths, the larger the median length of the paths.
+Furthermore, it is also expected that the higher the number of dominant paths, the larger the median length of the paths.
 Fig.
 6 (c) confirms this end by showing a strong linear relationship between them.
-Figure 5: (a) Dominant path length vs dominant path index for some illustrative measured channels; (b) scatter plot of the median value of the dominant path length vs the delay spread for the set of measured channels; (c) scatter plot of the median value of the dominant path length vs the number of dominant paths for the set of measured channels.
 IV-C Comparison with Other Works
 The MPM parameter distributions that have been just presented are herein compared with those heuristically or empirically obtained in previous scientific works.
-For example the presented statistical analysis confirms the assumptions made in [21], which statistically extended the MPM in (1) relying on the initial idea presented in [10].
-Hence, it was inferred in [21] that the path gains $g_{i}$ turn out to be log-normally distributed variables multiplied by a random flip sign.
-The main difference is found for the statistics of $N$, supposed to have a Poisson distribution in [21], but that results into a GEV one.
-This difference was already hypothesized in [21] and is also justified by the different considered databases.
-The path lengths, $d_{i}$, were assumed to be log-normally distributed in [21], while the mixture of a Weibull and GEV has been obtained in this work.
+For example the presented statistical analysis confirms the assumptions made in [Tonello12], which statistically extended the MPM in (1) relying on the initial idea presented in [Tonello07bis].
+Hence, it was inferred in [Tonello12] that the path gains $g_{i}$ turn out to be lognormally distributed variables multiplied by a random flip sign.
+The main difference is found for the statistics of $N$, supposed to have a Poisson distribution in [Tonello12], but that results into a GEV one.
+This difference was already hypothesized in [Tonello12] and is also justified by the different considered databases.
+The path lengths, $d_{i}$, were assumed to be lognormally distributed in [Tonello12], while the mixture of a Weibull and GEV has been obtained in this work.
 Additionally, the characterization method proposed in this paper provides a statistics for the other parameters, i.e., $a_{0}$, $a_{1}$, and $A$.
-Other results concerning the MPM parameters distribution are listed in [34](Tab.
-1), that extends the model in (1) to the MIMO transmission.
-In this case, the path gains $g_{i}$ are assumed uniformly distributed in $[-1,1]$ and the attenuation coefficient $a_{1}$ turns out to be constant.
+Other results concerning the MPM parameters distribution are listed in [Hashmat2011, Table 1], that extends the model in (1) to the MIMO transmission.
+In this case, the path gains $g_{i}$ are assumed uniformly distributed in [1,1][-1,1] and the attenuation coefficient $a_{1}$ turns out to be constant.
 While, $K$ is normally distributed and $a_{0}$ exhibits a shifted exponential distribution.
 Instead, the channel median $A$ is uniformly or exponentially distributed depending on the considered channels.
-The same results are also summarized in [35, Tab.
-2.10].
+The same results are also summarized in [Lampe16, Table 2.10].
 IV-D Relationship Between Parameters
 The development of MPM-based models requires a complete characterization of the involved parameters.
-This includes both their marginal PDFs, which have given in the previous subsection, and their joint distribution, which are examined in this one.
+This includes both their marginal PDFs, which have given in the previous subsection, and their conditional distribution, which are examined in this one.
 First, the relation between the path length, $d_{i}$, and the number of dominant paths, $N$, is explored.
-To this end, the probability distribution of the path length conditioned on $N$, $P_{r}\left(d_{i}|N\right)$ is estimated.
+To this end, the probability distribution of the path length conditioned on $N$, $f\left(d_{i}|N\right)$ is estimated.
 Obtained results are given in Fig.
-6, which gives additional details on the already observed bimodal behavior of the EPDF of $d_{i}$.
-As seen, there are almost no paths with lengths in the range from $1000-3000$ m, approximately, except for the case where $N$ is very large, in which some paths with lengths in this range can be found.
+7, which gives additional details on the already observed bimodal behavior of the EPDF of $d_{i}$.
+As seen, there are almost no paths with lengths in the range from 100030001000-3000 m, approximately, except for the case where $N$ is very large, in which some paths with lengths in this range can be found.
 The values of $d_{i}$ are concentrated in the low length region and in the region around 3000 m, being (approximately) uniformly distributed from zero (or almost zero) in the low length region.
@@
 Figure 7: Conditional probability of the path length, $d_{i}$, given the number of dominant paths, $N$, estimated over the set of measured channels.
-Next, the relation between the absolute value of the path gain and the path length is assessed by estimating the conditional PDF of $|g_{i}|$ given $d_{i}$, $f_{|g_{i}|}\left(|g_{i}|\big{|}d_{i}\right)$.
+Next, the relation between the absolute value of the path gain and the path length is assessed by estimating the conditional PDF of $|g_{i}|$ given $d_{i}$, $f\left(|g_{i}|\big|d_{i}\right)$.
 Results are displayed in Fig.
-7.
+8.
 Interestingly, $|g_{i}|$ tends to a uniform distribution in the interval $(0,1)$ for very low values of $d_{i}$.
 As the path length increases, the range of values of $|g_{i}|$ decreases until it tends, again, to a uniform one in the interval $(0,1)$ for path lengths around $3000$ m.
-There is a clearly noticeable gap for path lengths between $1000-2500$ (approximately), where the few existing paths have very low gains.
+There is a clearly noticeable gap for path lengths between 100025001000-2500 (approximately), where the few existing paths have very low gains.
 Figure 8: Conditional PDF of $|g_{i}|$ given the path length, $d_{i}$, estimated over the set of measured channels.
 Finally, the correlation between the path gains is explored.
-Since the decimation process described in Section III.III-A yields a different number of paths for each measured channel, the following procedure is employed to allow comparing results from different channels.
+Since the decimation process described in Section III-A yields a different number of paths for each measured channel, the following procedure is employed to allow comparing results from different channels.
 In order to distinguish the MPM model parameters before the decimation process is applied from the ones resulting after it, let us denote the latter by using tilded symbols.
-Hence, the original set of path lengths and path gains are expressed as $d_{n}$ and $g_{n}$, respectively, with $10\leq n\leq N-$ 1, where $N=$2554, and their corresponding sets after the decimation process as $\tilde{d}_{i}$ and $\tilde{g}_{i}$, respectively, with $0\leq i\leq\tilde{N}$, where $\tilde{N}$ is the number of dominant paths after the decimation process.
+Hence, the original set of path lengths and path gains are expressed as $d_{n}$ and $g_{n}$, respectively, with $0\leq n\leq N-1$, where N=N=2554, and their corresponding sets after the decimation process as $\tilde{d}_{i}$ and $\tilde{g}_{i}$, respectively, with $0\leq i\leq\tilde{N}$, where $\tilde{N}$ is the number of dominant paths after the decimation process.
 These magnitudes are related as
@@
 As seen, the ACF decreases very rapidly, with values below 0.5 for $n\geq 8$ and below 0.3 for $n\geq 40$, and slightly regrows for large values of $n$ due to the behavior already shown in Fig.
-7.
+8.
 Figure 9: Average and envelope of the normalized deterministic ACF of $|g_{n}|$ computed over the set measured channels.
-V Conclusion
+V Applicability of the study and influence on the performance of PLC systems
+This section discusses the use of the statistical analysis of the MPM parameters given in this work to the development of a stochastic MPM-based channel generator and also explores the relation between the model parameters and the performance of indoor PLC systems.
+V-A Application to the development of a stochastic MPM
+While the development of a statistical MPM-based channel generator is beyond the scope of this work, information provided herein can be used as a starting point to generate realizations of CFRs according to the MPM.
+To this end, the steps given below can be followed:
+  1. 1.
+    Generate a realistic value of the average channel gain, $G\;\textrm{(dB)}$}.
+To this end, the EPDF of $G\;\textrm{(dB)}$} for the set of measured channels is given in Fig.
+10 (a), along with the PDF of the distribution that gives the best fit, whose parameters are provided in Table VI.
+  2. 2.
+    Generate a realistic value of the delay spread.
+This can be obtained by leveraging the relation between the average channel gain and the delay spread depicted in Fig.
+10 (b), where the logarithm is taken to avoid the heteroscedasticity in the regression line, which is given by
+    $\ln(\widehat{\textrm{ds}}(\mu\textrm{s}))=\alpha+\beta\cdot G(\textrm{dB}),$ (17)
+    with $\alpha=-1.7499$ and $\beta=-2.7630\cdot 10^{-2}$. The distribution of the residuals of the fitting in (17) is given in Table VI.
+  3. 3.
+    Generate the values of $a_{0}$, $N$ and $A$ using the expressions given in Table III.
+  4. 4.
+    Since $a_{1}$ appears to be uncorrelated with the remaining parameters of the model, it can be generated independently according to the distribution given in Table V.
+Sections IV-B and IV-D have revealed useful information about the probability distributions of $d_{i}$ and $g_{i}$ and their relation to other parameters of the MPM.
+However, further research is required to generate realistic sequences of these magnitudes.
+Our preliminary analysis suggests that the distribution of $d_{i}$ conditioned on $N$ can be modelled as a mixture of the Pearson and GEV distributions.
+Nevertheless, a larger number of measured channels are required to obtain accurate estimates of the distribution parameters for high values of $N$.
+A mathematical fitting of the conditional PDF of $|g_{i}|$ given $d_{i}$, illustrated in Fig.
+8, is also required.
+The generation of sequences of $d_{i}$ requires knowledge of their joint distributions, which seems to be unrealistically complex.
+However, at least their covariance is required to obtain a linear model.
+Figure 10: (a) Empirical and fitted PDFs of the average channel gain; (b) scatter plot of the relation between the logarithm of the delay spread and the average channel gain.
+Table VI: Fitting distributions of the average channel gain and of the residuals of the fitting in (17)
+Channel
+parameter
+Distribution
+Distribution
+parameters
+Average channel gain
+$G$ (dB)
+GEV
+k=0.2984k=-0.2984
+$\sigma=13.9104$
+$\mu=-43.025$
+Residuals of the fitting in (17)
+Normal
+$\mu=1.2854$
+$\sigma=0.3328$
+V-B Relation between the MPM parameters and the performance of indoor PLC systems
+[…]
+For the sake of simplicity, continuous OFDM symbol transmission is assumed.
+A noise model with Gaussian distribution and PSD given by $S_{N}(f)=a+bf^{c}$ (dBm/kHz), with $f$ in MHz, $a=-1.7542\cdot 10^{3}$}, $b=1.6594\cdot 10^{3}$} and $c=-4.1565\cdot 10^{-3}$} is used [Cortes23].
+Fig.
+11 depicts the scatter plot of the PHY bit rate versus three key parameters of the MPM: the effective number of paths, $N$; the attenuation coefficient, $a_{0}$; and the normalization coefficient, $A$.
+As expected, performance decreases as $N$ and $a_{0}$ increases, in the former case because of the increased frequency selectivity of the channel and the latter because of the increased attenuation.
+As shown in Table VII, in both cases the relation can be fitted by means of a third-order polynomial function, although the cubic and quadratic terms have lower influence in the case of $a_{0}$, where the relation is almost linear.
+Regarding the relation to the normalization coefficient displayed in Fig.
+11 (c), the performance increases with $A$.
+This is coherent with the fact that $A=1/\underset{i}{\max}(|g_{i}|)$, and the greater the absolute value of the path gains, the larger the frequency selectivity of the channel response.
+Figure 11: Scatter plot and fitting curve of the estimated PHY bit rate (Mbit/s) versus (a) the number of effective paths, $N$; (b) the attenuation coefficient $a_{0}$; (c) the normalization coefficient, $A$.
+Table VII: Parameters of the fitting expressions displayed in Fig. 11.
+Fitting expression $\bm{\alpha}$ $\bm{\beta}$ $\bm{\gamma}$ $\bm{\delta}$
+$\widehat{R}\textrm{ (Mbit/s)}=\alpha+\beta\cdot N+\gamma\cdot N^{2}+\delta\cdot N^{3}$ $8.1687$ 2.7978-2.7978 $3.8227\cdot 10^{-3}$ $1.8133\cdot 10^{-6}$
+$\widehat{R}\textrm{ (Mbit/s)}=\alpha+\beta\cdot a_{0}+\gamma\cdot a_{0}^{2}+\delta\cdot a_{0}^{3}$ $7.0758\cdot 10^{2}$ $1.5554\cdot 10^{5}$ $1.359\cdot 10^{8}$ $3.5222\cdot 10^{10}$
+$\widehat{R}\textrm{ (Mbit/s)}=\alpha\cdot e^{\beta\cdot A}+\gamma$ $4.8969\cdot 10^{-2}$ 21.0197-21.0197 $6.9578\cdot 10^{2}$ -
+VI Conclusion
 The MPM is one of the most common top-down models adopted in the literature about PLC.
-This paper explores the parameterization of this model, which was originally proposed for low-voltage power distribution networks in the band up to 20 MHz, for the indoor broadband scenario in the band up to 80 MHz.
+This paper explores the parameterization of this model, which in the PLC context was firstly applied to low-voltage power distribution networks in the band up to 20 MHz, for the indoor broadband scenario in the band up to 80 MHz.
 Since the number of parameters that result from the extension of the MPM model to this context is extremely large, some of them have been imposed on the basis of common sense intuition, others, such as the attenuation parameters and the set of initial path lengths, have been chosen according to physical assumptions, and the remaining ones are analytically retrieved by means of a WLS fitting.
@@
 Suitable distributions that fit the empirical PDF of each parameter are also given.
-Finally, the relation between the most relevant parameters of the model is explored.
-Provided results will aid the development improved channels models for indoor broadband PLC.
+The relation between the most relevant parameters of the model is explored.
+Finally, the applicability of the provided results for the development of stochastic MPM models for indoor broadband PLC and the influence of the MPM parameters on the performance of PLC system are discussed.